FA2: A metal cylinder has a mass of 100.0 grams. If it measures 4.0 in high and 1.50 in wide, what is its density (ρ) in g/cm3 if 1.0 in=2.54 cm?
1
Expert's answer
2020-10-05T10:54:11-0400

"4.0in=10.16cm"


"1.50in=3.81cm"


"V=\frac{\pi d^2}{4}h=\frac{3.14\cdot 3.81^2}{4}\cdot10.16\approx116(cm^3)"


"\rho=m/V=100.0/116=0.862(g/cm^3)"